Consider the following simplified balance sheet for a bank.  Suppose the required reserve ratio decreases from 10 percent to 5 percent. By how much will bank deposits increase?   Provide your answer in dollars measured in thousands rounded to two decimal places. Do not include any symbols, such as "$," "=," "%," or "," in your answer.

Assets

Liabilities

Reserves

$10 thousand

Deposits

$100 thousand

Consider the following simplified balance sheet for a bank.  Suppose the required reserve ratio decreases from 10 percent to 6 percent.  By how much will the supply of money increase?   Provide your answer in dollars measured in thousands rounded to two decimal places. Do not include any symbols, such as "$," "=," "%," or "," in your answer.

Assets

Liabilities

Reserves

$10 thousand

Deposits

$100 thousand

Suppose that real GDP is currently $18.2 trillion, potential GDP is $23.7 trillion, the government purchases multiplier is 1.1, and the tax multiplier is -1.1. Holding other factors (such as prices and interest rates) constant, how will taxes (T) need to change to bring the economy to equilibrium at potential GDP?  Provide your answer in dollars measured in trillions rounded to two decimal places. Use a negative sign "-" for negative changes. Do not include any symbols, such as "$," "=," "%," or "," in your answer.
Answer #1

solution:

if required reserve ratio decreases from 10 percent to 5 percent. By how much will bank deposits increase:

Answer is $5 thousand

If reserve ration is decreased to 5% from 10%, This $ 5 thousand ( $100 thousand x 10% - $ 100 thousand x 5 % )

calculation:

When reserve ratio is 10% or 0.1

then bank's reserve is,

Reserve = Deposit * Required reserve ratio = 100 * 0.1 = 10 thousand dollar.

So, after keeping 10% reserve the bank's deposit amount is (100 - 10) = 90 thousand dollar.

When reserve ratio is 5% or 0.05

then bank's reserve is,

Reserve = Deposit * Required reserve ratio = 100 * 0.05 = 5 thousand dollar.

So, after keeping 5% reserve the bank's deposit amount is (100 - 5) = 95 thousand dollar.

Therefore, when reserve ratio falls from 10% to 5% then the bank's deposit increases by

(95 - 90) = 5 thousand dollar.

if required reserve ratio decreases from 10 percent to 6 percent.  By how much will the supply of money increase is:

Answer is $ 1 thousand

If the reserve ratio is decreased to 6 % from 10%,

The supply money increased by is:

required reserve ratio = 10%

deposit ratio = 100 thousands

RR = 10 thousand

ER = 0

rr = 7%

D = 100 thousand

RR = 6 thousand

ER = 4 thousand

money multiplier = 1/rr = 1/0.06

supply of money = 4000 * 1/0.06

= 66666.67

The supply money increased by = 66.67 thousand